THE DISPLACEMENT OF A PARTICLE STARTING from rest and moving under a constant acceleration is calculated by the formula S= 1/2 at^2 .if there occurs an error of 30% in the measurement of time interval , what error will be introduced in the calculation of S?
ANS 69%

Grade:12

3 Answers

RAJEEV CHOUDHARY
10 years ago
why 69%.if we take 30% less i.e.0.7t then it comes out to be 51%
ria
10 years ago
that is the reason y i posted this question
Harsh davda
10 years ago
its simple.... error for time is dt/t=30/100=0.3 (here d means delta) that means dt=0.3t now S=1/2 at^2 after applying error it becomes.. S+dS =1/2a(t+dt)^2 =1/2a(t^2+2tdt+[dt]^2) ...... now S=1/2 at^2... hence they get cancelled... we r left with dS=atdt+a([dt]^2)/2 ... now put dt =0.3t you would get dS= 0.345at^2... you want dS/S.. divide the equation by S= 1/2 at^2 .. you would get dS/S= 0.69 =69%